How many bit strings can be formed by using seven 1s and nine 0s?
1
Expert's answer
2020-11-22T18:42:25-0500

There are Seven 1's and nine 0's so in all there are total 7 + 9 = 16 positions for the placement of 1's and 0's


Total number of the formation of bit strings is given by = positioning seven 1's in 16 positions and filling up the remaining 9 positions with 0's = 16C7 * 9C9 = 11440*1 = 11440


Hence, 11440 different bit strings can be formed by using seven 1s and nine 0s